# FRABench and GenEval: Scaling Fine-Grained Aspect Evaluation across Tasks, Modalities

### v2

@@ Appendix F The Prompt Template Used by UAs and TAs
Table 62: The example prompt templates for TAs in IG task.
@@
[Response 1]: [Text 1]: {response_1_text_1} [Text 2]: {response_1_text_2} …
[Response 2]: [Text 1]: {response_2_text_1} [Text 2]: {response_2_text_2} …
@@
Table 63: The example prompt templates for TAs in ITIG task with input.
@@
[Response 1]: [Text 1]: {response_1_text_1} [Text 2]: {response_1_text_2} …
[Response 2]: [Text 1]: {response_2_text_1} [Text 2]: {response_2_text_2} …
@@
Table 64: The example prompt templates for TAs in ITIG task without input.
Appendix G More Qualitative Examples
Table 65, Table 66, Table 67, Table 68 to 70 show four good qualitative examples of feedback generated by GenEval, corresponding to the NLG, IU, IG, and ITIG tasks, respectively. Table 71, Table 72, and Table 73 shows a relatively inferior example of feedback generated by GenEval.
Example 2. GenEval in Judging Responses on the Title Generation Sub-Task.
@@ Appendix G More Qualitative Examples
GenEval [Feedback]: Both responses provide appealing titles for blog posts about cabin rentals in Gatlinburg, but Response 2 stands out slightly more in terms of attractiveness. Response 1 offers titles that are straightforward and descriptive, but they are somewhat generic and lack a unique selling point. Response 2, on the other hand, uses more evocative language and phrases like "Discover the Serenity," "Gatlinburgś Best-Kept Secrets," and "Adventure Awaits," which are more likely to capture the readerś attention and entice them to read further. The titles in Response 2 are more likely to spark interest and curiosity, making them more appealing overall. [Result]: 4 5
@@
Table 65: An example of comparing the feedback generated by GenEval, Themis and GPT-4o in title generation sub-tasks of NLG. GenEval enables more granular evaluation compared to Themis.
Example 3. GenEval in Judging Responses on the Detailed Image Captioning Sub-Task.
@@
GenEval [Feedback]: Both responses accurately describe the objects and activities present in the image, focusing on the people and their respective activities on the water. Response 1 mentions the person standing on a paddleboard and another person sitting or kneeling on the same paddleboard, as well as the orange pedal boat with two people in it. Response 2 similarly describes the person kneeling on a paddleboard and another person standing and paddling, along with two individuals in a bright orange, inflatable boat. Both responses align well with the objects and activities depicted in the image. [Result]: 5 5
@@
Table 66: An example of comparing the feedback generated by GenEval, LLaVA-Critic, and GPT-4o in image caption generation sub-tasks of IU.
Example 3. GenEval in Judging Responses on the Image Generation Sub-Task.
@@
GenEval [Feedback]: The first image accurately depicts two birds chasing each other in the air, with the bird flying higher having a long tail and the other bird having a short tail. The characteristics of the birds, including their relative positions and tail lengths, are clearly represented. The second image, while showing two birds in flight, does not clearly depict the difference in tail length between the two birds. The tails appear to be of similar length, which does not align with the description provided. [Result]: 5 3
@@
Table 67: An example of comparing the feedback generated by GenEval, Claude-3.5 and GPT-4o in image generation sub-tasks of IG.
Example 4. GenEval in Judging Responses on the Storytelling Generation Sub-Task.
@@
[Uncaptioned image] Figure 19: 1. Oliver woke up one morning to find his shadow missing, only to discover it had come to life and was stealing other shadows around town. [Uncaptioned image] Figure 20: 2. Determined to set things right, Oliver embarked on an adventure, chasing his mischievous shadow through bustling streets and dark alleys. [Uncaptioned image] Figure 21: 3. Along the way, he encountered various scenes where shadows were missing, leaving people and objects looking strange and incomplete.
@@
Table 68: An example of comparing the feedback generated by GenEval, Qwen2VL-72B and GPT-4o in storytelling generation sub-tasks of ITIG. (Part 1)
Example 4. GenEval in Judging Responses on the Storytelling Generation Sub-Task.
@@
Table 69: An example of comparing the feedback generated by GenEval, Qwen2VL-72B and GPT-4o in storytelling generation sub-tasks of ITIG. (Part 2)
Example 4. GenEval in Judging Responses on the Storytelling Generation Sub-Task.
@@
GenEval [Feedback]: Based on the criterion of originality, Response 1 demonstrates a more coherent and imaginative narrative. It follows a clear storyline where Oliver discovers his shadow has come to life and goes on an adventure to reclaim it. The images and text are well-integrated, showing scenes of Oliver chasing his shadow and the moment he reclaims the stolen shadow. This response effectively captures the essence of the story and maintains a consistent theme throughout. Response 2, on the other hand, lacks coherence and originality. The narrative is disjointed and introduces a different character, Jack, and his imaginary friend, Mr. Shadow. The story does not align well with the task description and fails to create a clear and engaging storyline. The images and text do not complement each other effectively, and the narrative does not follow a logical progression. [Result]: 5 2
@@
Table 70: An example of comparing the feedback generated by GenEval, Qwen2VL-72B and GPT-4o in storytelling generation sub-tasks of ITIG. (Part 3)
1. In the following example of Harmfulness Judgment, we speculate that the misunderstanding of GenEval is caused by over-associating some specific image feature with the criterion, yet it might be usable in some circumstances.
Example 5. GenEval in Judging Responses on the Image Generation Sub-Task.
@@
GenEval [Feedback]: Based on the given criterion of harmfulness, the first image depicts a shadowy figure with an ominous appearance, which could potentially evoke discomfort or unease in some viewers. However, it does not explicitly promote, glorify, or depict violence, abuse, or harm. The second image shows a glowing light source with tree-like patterns in the background, which is neutral and does not evoke any sense of harm or distress. Therefore, the second image is less likely to cause psychological discomfort compared to the first. [Result]: 3 5
@@
Table 71: An example of comparing the feedback generated by GenEval, GPT-4o and Qwen2VL-72B in image generation sub-tasks of IG.
2. In the following example, all three models show poor alignment with human judgment, due to their general disability in image observation.
Example 6. GenEval in Judging Responses on the Image Generation Sub-Task.
@@
GenEval [Feedback]: The first image demonstrates higher fidelity to the description as it accurately portrays a cyber female with mechanical elements integrated into her eyes and lips, aligning with the specified features. The constellation background and stars are also well-represented, enhancing the alignment with the description. In contrast, the second image, while visually appealing, deviates slightly in fidelity. The mechanical features are less pronounced and do not clearly extend to the lips, and the background, though celestial, lacks the distinct constellation theme described. These discrepancies result in a lower fidelity score for the second image. [Result]: 5 3
@@
Claude-3.5 [Feedback]: Both images demonstrate a high level of fidelity to the given description of a "cyber female with mechanic eyes and lips, constellation background, stars." However, there are some differences in how closely they adhere to the specific details. Image 1 shows a single cyber female with mechanical elements incorporated into her design, particularly around her head. Her eyes appear to have a mechanical or enhanced quality, aligning well with the description. The lips are prominently red, which could be interpreted as a mechanical enhancement. The background clearly shows a starry, constellation-like setting, faithfully representing the described environment. Image 2 depicts two cyber females, which is not explicitly mentioned in the description. Both figures have mechanical elements, particularly in their attire and surrounding features. The eyes of both figures appear to have an enhanced, possibly mechanical quality. The lips are prominently red, similar to Image 1. The background also shows a starry sky with celestial bodies, adhering to the "constellation background, stars" part of the description. While both images capture the essence of the description, Image 1 adheres more closely to the specific details provided. It shows a single cyber female as described, whereas Image 2 presents two figures. Both images effectively represent the mechanical eyes, enhanced lips, and starry background. In terms of fidelity to the given description, Image 1 is slightly more accurate due to the depiction of a single figure as implied by the description. [Result]: 3 4
Table 72: An example of comparing the feedback generated by GenEval, GPT-4o and Claude-3.5 in image generation sub-tasks of IG.
3. In the following example, all three models come up with bad alignment with human judgment due to their poor math ability.
Example 7. GenEval in Judging Responses on the Data Analysis Sub-Task.
@@
Response 1: Person: George Washington Organization: troops Date: December 25, 1776 Action: crossed the Delaware River
@@
GenEval [Feedback]: The first image demonstrates higher fidelity to the description as it accurately portrays a cyber female with mechanical elements integrated into her eyes and lips, aligning with the specified features. The constellation background and stars are also well-represented, enhancing the alignment with the description. In contrast, the second image, while visually appealing, deviates slightly in fidelity. The mechanical features are less pronounced and do not clearly extend to the lips, and the background, though celestial, lacks the distinct constellation theme described. These discrepancies result in a lower fidelity score for the second image. [Result]: 5 3
@@
Table 73: An example of comparing the feedback generated by GenEval, GPT-4o and Qwen2VL-72B in data analysis sub-tasks of NLG.

### v3

@@ Appendix G Appendix G: Qualitative Examples
Table 58: The prompt templates for Task-specific Aspects in IG task.
You will be given an image description and two images generated by two models based on the image description. You are tasked with analyzing why, when evaluating the two images based on the given criterion, the evaluation result of the first image is {first_rating}, while the evaluation result of the second image is {second_rating}. {compare_description} Here is the data:
[BEGIN DATA]
[Criterion]: {criterion_description}
[Image Description]: {image_description}
[END DATA]
Here are the instructions to assess and compare the two images:
1. Carefully review every detail of the two images and image description, and provide detailed feedback analyzing why the first image {compare_word} the second in terms of evaluation results, based on the given evaluation criterion. Do not evaluate them in general terms or consider factors unrelated to the specified criteria, such as clarity or detail.
2. Based on the elements described in the image description, search for the required characteristics of each element, then compare them with the elements presented in the image to determine if they match, in order to perform the {criterion_name} evaluation.
3. Do not list points. Write a feedback paragraph of 50-100 words. After writing your feedback, assign two integer scores based on the given criterion (ranging from 1 to 5, with higher scores indicating better performance).
Table 59: The prompt templates for GPT-4o to generate feedback for human-annotated scores.
[BEGIN DATA]
[Criterion]: {criterion_description}
[Response 1]: {response_1}
[Response 2]: {response_2}
[END DATA]
Here are the instructions to assess and compare the two responses:
Table 60: The prompt templates for Universal Aspects only include the text output modality.
[BEGIN DATA]
[Criterion]: {criterion_description}
[END DATA]
Here are the instructions to assess and compare the two images:
Table 61: The prompt templates for Universal Aspects only include the image output modality.
@@
[Response 1]: [Text 1]: {response_1_text_1} [Text 2]: {response_1_text_2}
[Response 2]: [Text 1]: {response_2_text_1} [Text 2]: {response_2_text_2}
@@
Table 62: The example prompt templates for Task-specific Aspects in ITIG task with input.
@@
[Response 1]: [Text 1]: {response_1_text_1} [Text 2]: {response_1_text_2}
[Response 2]: [Text 1]: {response_2_text_1} [Text 2]: {response_2_text_2}
@@
Table 63: The example prompt templates for Task-specific Aspects in ITIG task without input.
Example 1. UFEval in Judging Responses on the Title Generation Sub-Task.
@@
UFEval [Feedback]: Both responses provide appealing titles for blog posts about cabin rentals in Gatlinburg, but Response 2 stands out slightly more in terms of attractiveness. Response 1 offers titles that are straightforward and descriptive, but they are somewhat generic and lack a unique selling point. Response 2, on the other hand, uses more evocative language and phrases like "Discover the Serenity," "Gatlinburgś Best-Kept Secrets," and "Adventure Awaits," which are more likely to capture the readerś attention and entice them to read further. The titles in Response 2 are more likely to spark interest and curiosity, making them more appealing overall. [Result]: 4 5
@@
Table 64: An example of comparing the feedback generated by UFEval, Themis and GPT-4o in title generation sub-tasks of NLG. UFEval enables more granular evaluation compared to Themis.
Example 2. UFEval in Judging Responses on the Detailed Image Captioning Sub-Task.
@@
UFEval [Feedback]: Both responses accurately describe the objects and activities present in the image, focusing on the people and their respective activities on the water. Response 1 mentions the person standing on a paddleboard and another person sitting or kneeling on the same paddleboard, as well as the orange pedal boat with two people in it. Response 2 similarly describes the person kneeling on a paddleboard and another person standing and paddling, along with two individuals in a bright orange, inflatable boat. Both responses align well with the objects and activities depicted in the image. [Result]: 5 5
@@
Table 65: An example of comparing the feedback generated by UFEval, LLaVA-Critic, and GPT-4o in image caption generation sub-tasks of IU.
Example 3. UFEval in Judging Responses on the Image Generation Sub-Task.
@@
UFEval [Feedback]: The first image accurately depicts two birds chasing each other in the air, with the bird flying higher having a long tail and the other bird having a short tail. The characteristics of the birds, including their relative positions and tail lengths, are clearly represented. The second image, while showing two birds in flight, does not clearly depict the difference in tail length between the two birds. The tails appear to be of similar length, which does not align with the description provided. [Result]: 5 3
@@
Table 66: An example of comparing the feedback generated by UFEval, Claude-3.5 and GPT-4o in image generation sub-tasks of IG.
Example 4. UFEval in Judging Responses on the Storytelling Generation Sub-Task.
@@
[Uncaptioned image] Figure 18: 1. Oliver woke up one morning to find his shadow missing, only to discover it had come to life and was stealing other shadows around town. [Uncaptioned image] Figure 19: 2. Determined to set things right, Oliver embarked on an adventure, chasing his mischievous shadow through bustling streets and alleys. [Uncaptioned image] Figure 20: 3. Along the way, he encountered various scenes where shadows were missing, leaving people and objects looking strange and incomplete.
@@
Table 67: An example of comparing the feedback generated by UFEval, Qwen2VL-72B and GPT-4o in storytelling generation sub-tasks of ITIG. (Part 1)
Example 4. UFEval in Judging Responses on the Storytelling Generation Sub-Task.
@@
Table 68: An example of comparing the feedback generated by UFEval, Qwen2VL-72B and GPT-4o in storytelling generation sub-tasks of ITIG. (Part 2)
Example 4. UFEval in Judging Responses on the Storytelling Generation Sub-Task.
@@
UFEval [Feedback]: Based on the criterion of originality, Response 1 demonstrates a more coherent and imaginative narrative. It follows a clear storyline where Oliver discovers his shadow has come to life and goes on an adventure to reclaim it. The images and text are well-integrated, showing scenes of Oliver chasing his shadow and the moment he reclaims the stolen shadow. This response effectively captures the essence of the story and maintains a consistent theme throughout. Response 2, on the other hand, lacks coherence and originality. The narrative is disjointed and introduces a different character, Jack, and his imaginary friend, Mr. Shadow. The story does not align well with the task description and fails to create a clear and engaging storyline. The images and text do not complement each other effectively, and the narrative does not follow a logical progression. [Result]: 5 2
@@
Table 69: An example of comparing the feedback generated by UFEval, Qwen2VL-72B and GPT-4o in storytelling generation sub-tasks of ITIG. (Part 3)
Example 5. UFEval in Judging Responses on the Image Generation Sub-Task.
@@
UFEval [Feedback]: Based on the given criterion of harmfulness, the first image depicts a shadowy figure with an ominous appearance, which could potentially evoke discomfort or unease in some viewers. However, it does not explicitly promote, glorify, or depict violence, abuse, or harm. The second image shows a glowing light source with tree-like patterns in the background, which is neutral and does not evoke any sense of harm or distress. Therefore, the second image is less likely to cause psychological discomfort compared to the first. [Result]: 3 5
@@
Table 70: An example of comparing the feedback generated by UFEval, GPT-4o and Qwen2VL-72B in image generation sub-tasks of IG.
Example 6. UFEval in Judging Responses on the Image Generation Sub-Task.
@@
UFEval [Feedback]: The first image demonstrates higher fidelity to the description as it accurately portrays a cyber female with mechanical elements integrated into her eyes and lips, aligning with the specified features. The constellation background and stars are also well-represented, enhancing the alignment with the description. In contrast, the second image, while visually appealing, deviates slightly in fidelity. The mechanical features are less pronounced and do not clearly extend to the lips, and the background, though celestial, lacks the distinct constellation theme described. These discrepancies result in a lower fidelity score for the second image. [Result]: 5 3
@@
Claude-3.5 [Feedback]: Both images demonstrate a high level of fidelity to the given description of a "cyber female with mechanic eyes and lips, constellation background, stars." However, there are some differences in how closely they adhere to the specific details. Image 1 shows a single cyber female with mechanical elements incorporated into her design, particularly around her head. Her eyes appear to have a mechanical or enhanced quality, aligning well with the description. The lips are prominently red, which could be interpreted as a mechanical enhancement. The background clearly shows a starry, constellation-like setting, faithfully representing the described environment. mage 2 depicts two cyber females, which is not explicitly mentioned in the description. Both figures have mechanical elements, particularly in their attire and surrounding features. The eyes of both figures appear to have an enhanced, possibly mechanical quality. The lips are prominently red, similar to Image 1. The background also shows a starry sky with celestial bodies, adhering to the "constellation background, stars" part of the description. While both images capture the essence of the description, Image 1 adheres more closely to the specific details provided. It shows a single cyber female as described, whereas Image 2 presents two figures. Both images effectively represent the mechanical eyes, enhanced lips, and starry background. In terms of fidelity to the given description, Image 1 is slightly more accurate due to the depiction of a single figure as implied by the description. [Result]: 3 4
Table 71: An example of comparing the feedback generated by UFEval, GPT-4o and Claude-3.5 in image generation sub-tasks of IG.
Example 7. UFEval in Judging Responses on the Data Analysis Sub-Task.
@@
Response 1: Person: George Washington. Organization: troops. Date: December 25, 1776. Action: crossed the Delaware River
@@
UFEval [Feedback]: The first image demonstrates higher fidelity to the description as it accurately portrays a cyber female with mechanical elements integrated into her eyes and lips, aligning with the specified features. The constellation background and stars are also well-represented, enhancing the alignment with the description. In contrast, the second image, while visually appealing, deviates slightly in fidelity. The mechanical features are less pronounced and do not clearly extend to the lips, and the background, though celestial, lacks the distinct constellation theme described. These discrepancies result in a lower fidelity score for the second image. [Result]: 5 3
@@
Table 72: An example of comparing the feedback generated by UFEval, GPT-4o and Qwen2VL-72B in data analysis sub-tasks of NLG.